Activity Title: Mechanical Energy


Learning Target : To determine the mechanical energy of a system. Consider a simple pendulum with a bob mass 4.2 kg



Questions:


1. If the speed of the bob at the lowest point B is 0.84 m/s , What is its kinetic energy at this point?


2. If points A and C are the highest points reached by the bob, What is the potential energy if h= 0.036m at these points.?


3. What is the kinetic energy of the bob points A and C ?


4. What is the potential energy at point B.?


5. What is the total mechanical energy at point A.?


6. What is the total mechanical energy at point B.?

1
Expert's answer
2022-01-26T17:46:32-0500

Gives

M=4.2kg

Point B velocity

"V_B=0.84m\/sec"

Point A and C hiegh


h=0.036m

Part(a)

Kinetic energy of lowest point

"K_B=\frac{1}{2}mv^2"


"K_B=0.5\times4.2\times(0.84)^2=1.48176J"

Part(b )

Point (A) and Point (B) potential energy

"PE=mgh"

PE="4.2\times9.8\times0.036=1.48176J"

"PE_{A}=PE_{C}=1.48176J"

Part (c)

Kinetic energy of point ( A)and (c)

"KE_{A}=KE_{C}=0"

At Point( a) and point( c) kinetic energy equal zero

part(d)

Potential energy at point B

"PE_{B}=0J"

At Point (B) potential energy is zero

part(e) total machanical energy point (A)

(Machanical energy )"_A" =

"KE_A+PE_B"

Machanical energy at point (A) =0+mgh

ME=mgh

ME=1.48176J

Part(f)

Machanical energy at point (B)

"ME=KE_B+PE_B"

"ME=\frac{1}{2}mv^2+0"

Mechanical energy at point B

"ME_B=1.48176+0=1.48176J"